OpenStudy (anonymous):

Two angles of a triangle measure 37° and 46°. What is the measure of the third angle of the triangle? A. 44° B. 53° C. 97° D. 107°

OpenStudy (texaschic101):

The interior angles of a triangle always add up to 180 degrees. Does that help you ?

OpenStudy (anonymous):

a little

OpenStudy (anonymous):

so um is it 97

OpenStudy (texaschic101):

a + b + c = 180 37 + 46 + c = 180 83 + c = 180 c = 180 - 83 c = 97 yep..you got it

OpenStudy (anonymous):

thanks for the help

OpenStudy (texaschic101):

anytime :)